OpenStudy (anonymous):

change -54 degrees to radian measure in terms of pi

OpenStudy (solomonzelman):

For example, if I want to convert \(\color{#000000 }{ \displaystyle 40^\circ }\) to radians, then I divide by 180 and multiply times \(\large\color{#000000 }{ \pi }\). \(\large\color{#000000 }{ \displaystyle 40^\circ =40\times \pi \div 180 = 2\pi/9 \color{gray}{\rm~ rad} =0.69813 \color{gray}{\rm~ rad}}\)

OpenStudy (anonymous):

i got 3pi/10 as my answer ; n ;

OpenStudy (danjs):

-54 degrees, that is in quadrant 4

OpenStudy (danjs):

just a sign change

OpenStudy (danjs):

yeah just remember a full circle is 2pi or 360 2pi rad = 360 deg use that ratio to convert units

OpenStudy (danjs):

or pi/180 as above, same thing

OpenStudy (danjs):

or use any degree you want, just anything you know to get from rad to deg or back